### v1

# Enhancing selectivity using Wasserstein distance based reweighing

Abstract
Given two labeled data-sets $\mathcal{S}$ and $\mathcal{T}$, we design a simple and efficient greedy algorithm to reweigh the loss function such that the limiting distribution of the neural network weights that result from training on $\mathcal{S}$ approaches the limiting distribution that would have resulted by training on $\mathcal{T}$.
On the theoretical side, we prove that when the metric entropy of the input data-sets is bounded, our greedy algorithm outputs a close to optimal reweighing, i.e., the two invariant distributions of network weights will be provably close in total variation distance. Moreover, the algorithm is simple and scalable, and we prove bounds on the efficiency of the algorithm as well.
[…]
1 Introduction
Deep learning has found applications in diverse areas ranging from organic chemistry to computer generated art. Its applicability is limited by the availability of large amounts of labeled training data. A typical all-to-all neural net of width $n$ (say $n=10000$) and depth $d$ (say $d=10$) has $\Theta(n^{2}d)$ ($\simeq 10^{8}$) weight parameters, and although many neural nets work well despite being somewhat over-parameterized, they still need the number of training examples to be of a similar or only slightly smaller order of magnitude.
This reliance on large amounts of training data leads to difficulties in training when we have multiple objectives. For example, suppose we separately gather training data for a neural net to perform classification tasks into a set of classes A using labeled data-set $\mathcal{S}$, and into a set of classes B using labeled data-set $\mathcal{T}$. Then, unless $\mathcal{T}$ and $\mathcal{S}$ have many data-points in common, we will not have enough examples to train a neural net for the classification into classes of $A\times B$. In this paper, we address this problem by designing a scalable algorithm that reweighs dataset $\mathcal{S}$ (using $\mathcal{T}$), so that training on the reweighted $\mathcal{S}$ leads to network that is close to one obtained by training on $\mathcal{T}$. Moreover, we apply it to a drug discovery application and obtain wet-lab verified results.
Our main contribution is theoretical. Algorithm 1 deliberately introduces distribution skew and reweighs the labeled training data-set $\mathcal{S}$ using the data-set $\mathcal{T}$ so that if we train a neural network on the reweighed $\mathcal{S}$ for a long enough period of time then its weights will be "tilted" so that the classification error on classes $B$ in $\mathcal{T}$ will also be reduced. The amount of reduction is determined by the choice of tilt parameter $\alpha$ in Algorithm 1. Theorems 5.45.135.15, and 5.16 formally show correctness and efficiency of Algorithm 1. In particular, Theorem 5.4 proves correctness and also justifies the choice of Wasserstein metric in Algorithm 1. Theorems 5.135.15 and 5.16 provide efficiency guarantees for Algorithm 1.
It has been known since the 1980s that greedy algorithms, like Algorithm 1, have poor approximation guarantees for computing Wasserstein distance [20]. Therefore, we need to assume and exploit some property of about our input instances to get around the lower bound in [20]. This critical property turn out to be: input instances must admit a small size covering. This small covering assumption is used in both the random sampling and the greedy reweighing steps of our algorithm. The connection to random sampling is not surprising and it comes from known techniques – the union bound in the large deviations proof. However, the connection between small coverings and greedy matching algorithms (at its core computing Wasserstein distance is equivalent to computing minimum weight matchings) is somewhat surprising, since we are not aware of results obtaining sharper guarantees on approximate minimum weight matching, based on the covering properties of the input data-set.
Organization: In Section 2, we describe an example application to drug discovery. More theoretically inclined readers may skip this section and any references to it. In Section 3, we discuss prior work from the areas of machine learning theory, algorithms and computational drug discovery, relevant to our paper. In Section 4, we present Algorithm 1 and provide a technical overview of our paper – how our various theoretical results fit together. Section 5 presents the theorem statements. In particular, Theorem 5.4 explains why the Wasserstein metric is an intuitive and appropriate choice of metric for Algorithm 1; Theorems 5.135.15 and 5.16 show that the greedy random sampling based algorithm can compute the minimum weight bipartite matching, and hence Wasserstein distance, in near linear time; and they provide an upper-bound on the approximation error under our low metric entropy assumption. Thus showing that Algorithm 1 is scalable. Finally, the supplement contains missing proofs.
2 Example application: drug discovery
Multicriteria optimization problems involving data-sets that admit small coverings arise naturally in drug discovery (see Section 4 for details). Therefore, as a concrete motivating example, we illustrate an application of Algorithm 1 to a toy problem in this area.
In drug discovery, typically, one wants to isolate small molecules (inhibitors) that bind strongly to a given enzyme, but often we want to exclude small molecules that bind to another similar enzyme. For example, MNK1 and MNK2 are two structurally similar kinases (a kinase is an enzyme for phosphorylation or de-phosphorylation of proteins). We want to identify small molecules that bind strongly to MNK2 (MNK2 hits), but we also prefer that the identified small molecules not bind to MNK1 (MNK1 non-hits). In other words, we want to isolate molecules that are selective for MNK2 over MNK1.
In the in-silico experiments, we were able to increase the percentage of MNK1 non-hits in our set of top predicted MNK2 hits – the selectivity – from 54% to 95% on holdout data, using the reweighing procedure in Algorithm 1. We used a relatively small training set of about $250$K small molecules in total; labeled as MNK1 non-binders, and MNK2 binders and non-binders; and a small holdout set of $7$K small molecules that consists of molecules which are labeled as: MNK2 hits (binders), and MNK1 hits (binders) or MNK1 non-hits (non-binders). In Figure 1, for the neural network models with and without reweighing, we plot the cumulative number of MNK1 non-hits on the $y$-axis; and on the $x$-axis any given point, say $k$, represents the top $k$ predicted MNK2 hits from the examples in the holdout set.
Figure 1: Selectivity of reweighed (using Algorithm 1) and baseline (without reweighing) neural nets. Note that this increase in selectivity from 54% to 95% came without any significant change in the validation loss – the AUC for the classification of MNK2 binders vs non-binders remained around $0.6$ in both cases.
While we can not make our training data-sets and code public for proprietary reasons, we were able to experimentally (in wet-lab) verify that two out of the top fifty (actually 43, since 7 out of 50 molecules could not be synthesized and tested) predicted selective small molecules, obtained by running our neural network model on the Enamine 1.9B molecules catalog (https://enamine.net), were indeed selective for MNK2 over MNK1.111The compounds are Z1918489591 and Z5890616727 in the enamine catalog. That is a success rate of roughly 5% on this admittedly small sample set. We do note that the results are from a single point concentration assay and can be noisy.
We are not aware of other such multi-target prediction results in DNA encoded library (DEL) space (see [22] for background), where one simultaneously predicts hits/non-hits against two or more proteins. However, the success rates for single target experiments with traditional high-throughput screening is $\sim 1\%$ (see for example the discussion in [17]) and it is generally accepted that multi-target prediction is a harder problem.
More importantly, the two predicted and assay tested molecules provide a degree of verification for our experimental application (which has been the motivation, but is not the focus of this paper).
Figure 2: Two predicted and verified selective MNK1 non-hits and MNK2 hits from the Enamine catalog. The enzyme activity was found to be above 50% for MNK1 but below 50% for MNK2 at $10\mu$M concentration for each of the two small molecules: $\sim 20\%$ vs $70\%$ and $\sim 39\%$ vs $59\%$. Note that these values are from single point concentration assay and can be noisy.
3 Related work
The question of learning with differing test and train distributions has been well investigated in the machine learning community under different names: ditribution shift and covariate shift, see for example the book [19], the papers [24, 21, 10, 3, 13] and [7], to name just a few. The question is also relevant to our paper since our algorithm can be used to reweigh the train data-set to bring the post training neural network weights closer to what they would have been, had we trained based on the test distribution. However, prior results rely on estimating the train and test distributions. For example, parameter estimation of the densities followed by change of variable using the Jacobian. Assuming a logarithmic number of features, that leads to a $\tilde{O}(n^{2})$ algorithm for distribution skew correction ($n$ being the training and test data-set size) – much more efficient than Wasserstein distance computation that requires solving a $\Theta(n^{2})$ sized linear program. That explains why the Wasserstein distance based ideas are less explored in this context, so far. Distribution shift correction has also found new applications in domain adaptation literature. However, the only prior theoretical works involving Wasserstein distance computation that we found in this area were [8] and [16], which focus on exact solution of the Wasserstein distance problem.
In a high dimensional feature space, the density estimation can be inaccurate since the number of samples required increases exponentially in number of features for any formal guarantee (as can be seen from large deviation bounds [9]). Moreover, if we are only interested in partial tilting of one distribution towards another, as in Algorithm 1 (where $\alpha$ controls the amount of tilt), then it is reasonable to look for approximate but efficient computation of Wasserstein distance. That is what we do in this paper using Algorithm 1, which is a $\tilde{O}(n^{2})$ time algorithm as well. We now have the added advantage of a provable upper bound on the test-train prediction error, before and after tilting using Algorithm 1, from Theorem 5.4.
The problem of efficient Wasserstein distance computation has also received much attention in the algorithms community. The paper [23] studies the equivalence between Wasserstein distance computation and matching algorithms in the metric space setting. Efficient matching algorithms have been well studied in literature for five decades. The optimal algorithm for computing weighted matchings is due to Gabow and Tarjan [11] and runs in time $O(m\sqrt{n})$, where $m$ is the number of edges and $n$ the number of vertices in the graph. Since then more sophisticated algorithms have been designed, see for example  [25, 14, 23] and [1] to name a few. However, under our assumptions even the simple greedy algorithm performs remarkably well, and it scales efficiently for large training data-sets.
Reingold and Tarjan [20] showed that the greedy algorithm has an abysmal approximation ratio of $n^{\log_{2}3/2}$ for bipartite graphs. In this paper, we show in Theorem 5.13 that the approximation ratio of the greedy algorithm is much better under our bounded metric entropy assumption than the lower bound in [20]. Hence, an assumption about a covering property of the input leads to more optimal matchings – a somewhat surprising algorithmic result that may be of independent interest.
Finally, the idea of using deep learning for drug discovery has gained popularity in pharmaceutical research over the last few years, especially given the amount of data now available [18]. The paper [17] shows that neural nets can be trained on DNA encoded chemical libraries to identify new small molecules that bind to a given protein target. It is particularly relevant to this work, as we build upon that. Our work extends their work by allowing us to select molecules that bind to one protein target and not to another. Other papers in this rapidly growing area include [26, 15] and [12].
4 Problem statement and overview of results
Suppose we are given two labeled training data sets, say $\mathcal{S}$ and $\mathcal{T}$ for two different classification tasks. Moreover, let’s assume that the points in the data-sets are weighted according to two different probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ respectively. Our goal is to reweigh $\mathcal{S}$, i.e., "tilt" $\mathbb{P}_{\mathcal{S}}$ towards $\mathbb{P}_{\mathcal{T}}$, and train a neural net classifier so that the limiting distribution of network weights is closer to the one that would be obtained from training using $\mathbb{P}_{\mathcal{T}}$. We assume labels of $\mathcal{S}$ are known, and the labels of $\mathcal{T}$ may be unknown or they may be known but $|\mathcal{T}\cap\mathcal{S}|$ may be small. Our reweighing algorithm handles both cases. Furthermore, our reweighing procedure (Algorithm 1) works efficiently on very large data-sets, with provable guarantees.
For our drug discovery example, the set $\mathcal{T}$ consists of a subset of small molecules [^2] As an aside, each small molecule is usually mapped to a 2K character long binary string (fingerprint) of features. Thus, in this context, one may think of the underlying space of small molecules as a subset of the boolean hypercube in dimension 2K. labeled non-binders (non-hits) for the protein MNK1, and the full labeled training set consists of molecules that are binders and non-binders for the protein MNK2, while the set $\mathcal{S}$ is not the full training set but just the set of binders to MNK2333Note that we have the labeled set of non-binders to MNK2 in our training examples as well, but they remain unaffected by the reweighing, though they are used in training as well. . Here the labels of the molecules in $\mathcal{T}$ are known but not necessarily on the same molecules as $\mathcal{S}$. $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ may be assumed to be uniform distributions supported on $\mathcal{S}$ and $\mathcal{T}$ respectively. Given a new small molecule, one now wants to compute the likelihood that it is a binder for MNK2 and a non-binder for MNK1. Such models can allow us to make predictions on large commercially available catalogs and enrich compounds that have high likelihood to bind to MNK2 but not MNK1. This formulates the drug discovery application in our formal notation.
The rest of the paper concentrates on providing a theoretical explanation for why Algorithm 1 should work as intended and scale well in general; beyond the specific experiments with MNK1-MNK2.
Algorithm 1 Reweigh Distribution and Train
1:Input: Two data-sets: $\mathcal{S}$ and $\mathcal{T}$ of size $n$ each, points weighed according to $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ respectively, and a tilt factor $\alpha\in[0,1]$.
2:Output: Compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ on $\mathcal{S}$ such that the invariant distribution of network weights of a neural net model, trained using SGD with dataset $\mathcal{S}$ and weights $\mathbb{P}_{\mathcal{S}}^{\prime}$, will be closer (in Wasserstein metric) to the invariant distribution of network weights of a neural net model trained using SGD on $\mathcal{T}$ with points weighted as $\mathbb{P}_{\mathcal{T}}$.
3$\triangleright$ RandomSample returns an empirical probability distribution computed from sample size $m$.
4$\triangleright$ $R_{\mathcal{S}}\subseteq\mathcal{S}$ and $R_{\mathcal{T}}\subseteq\mathcal{T}$ denote the random sample of points from their respective ground sets.
5:$\mathbb{P}_{R_{\mathcal{S}}}:=\mathrm{RandomSample}_{m}(\mathcal{S},\mathbb{P}_{\mathcal{S}})$
6:$\mathbb{P}_{R_{\mathcal{T}}}:=\mathrm{RandomSample}_{m}(\mathcal{T},\mathbb{P}_{\mathcal{T}})$
7$\triangleright$ Obtain a $\alpha$-tilted version of $\mathbb{P}_{R_{\mathcal{S}}}$ that’s close to $\mathbb{P}_{R_{\mathcal{T}}}$ using greedy minimum weight metric bipartite matching algorithm (ScaledGreedyReweight in supplement)
8:$\mathbb{P}_{R_{\mathcal{S}}}^{\prime}:=\mathrm{ScaledGreedyReweight}(\mathbb{P}_{R_{\mathcal{S}}},\mathbb{P}_{R_{\mathcal{T}}},\alpha)$
9$\triangleright$ Obtain a reweighted version of $\mathcal{S}$
10:$\mathbb{P}_{\mathcal{S}}^{\prime}=(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{R_{\mathcal{S}}}^{\prime}$.
11$\triangleright$ Train neural net on $\mathbb{P}_{\mathcal{S}}^{\prime}$.
12:Use stochastic gradient descent (SGD) to train the neural net using $\mathbb{P}_{\mathcal{S}}^{\prime}$.
A difference in $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ results in a difference in the convergence point of the weights in any neural net training procedure, like SGD. This is because the loss functions in the SGD algorithm will differ in their weights, though they may have the same form. Therefore, given two mean squared error loss functions weighted with different probability distributions, say $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, on each of their terms, a natural question is: what is the relation between the output of two neural nets that are trained using the two different loss functions?
The expected output of any neural net, on any given input, at the end of a long enough training period, depends on the invariant measure of the weights from the SGD training algorithm. The $1$-Wasserstein distance [^4] See, for example [4] and the references therein for background on the Wasserstein distances. between the weights in the loss functions is the same as the $1$-Wasserstein distance between the occurrence frequencies of data-points in the two underlying data-sets $\mathcal{S}$ and $\mathcal{T}$. Our first theoretical contribution, Theorem 5.4, shows that $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, the 1-Wasserstein distance between the loss function weight distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$, upper-bounds the total variation distance between the invariant measures of two such neural nets. Therefore, if the data-sets $\mathcal{S}$ and $\mathcal{T}$ are not identical, i.e., $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ do not have the same support set, then the best we can hope for is a small $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, in order to obtain similar neural net models for interchangeability. The above explains our choice of the $1$-Wassertein metric in Algorithm 1 (see also Remark 5.5 for comparing with the Levy metric).
Let’s assume $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ is large, so that we need to reweigh $\mathcal{S}$. For reweighing, we do not remove any examples from $\mathcal{S}$, instead we prefer to increase the weight of some already present examples in $\mathcal{S}$. Therefore, we compute a distribution $\mathbb{P}_{\mathcal{S}}^{\prime}$ with the same set of support as $\mathbb{P}_{\mathcal{S}}$, such that it minimizes $1$-Wasserstein distance between $(1-\alpha)\mathbb{P}_{\mathcal{S}}+\alpha\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}^{\prime}$, for some fixed choice of tilt factor $\alpha\in[0,1]$.555 The optimum value of $\alpha$ can be chosen by binary search after running multiple training evaluations. We use $\mathbb{P}_{\mathcal{S}}^{\prime}$ as the new set of weights for neural net training. Note that, for $\alpha$ close to $1$, the distribution will be closest to $\mathbb{P}_{\mathcal{T}}$ while being supported on $\mathcal{S}$.
Note that the optimal $\mathbb{P}_{S}^{\prime}$ mentioned above can be computed by solving a linear program that closely resembles the $1$-Wasserstein distance computation linear program. However, the size of the linear program would be quadratic in the size of the data-sets, making the computation intractable for most practical data-sets. [^6] A typical large data-set has 10-100M examples, and computing $W_{1}$ over two such data-sets requires solving a linear program – a $\Theta(n^{3})$ time procedure, resulting in the order of $10^{24}$ computational operations! Therefore, we look for inaccurate but efficient algorithms and a natural candidate is the randomized greedy algorithm below.
The 1-Wasserstein metric has an equivalent interpretation as a optimal transport problem. See Theorem 5.6 (essentially repeated from [23]) for a formal statement that reduces it to the metric minimum weight bipartite matching problem [20].
Given a bipartite graph with vertices embedded in a metric space, the metric minimum weight bipartite matching problem asks to compute a minimum weight matching, where the weight of a matching is the sum of the lengths of edges in the matching.
One tractable way to compute a minimum weight bipartite matching is to use a faster but sub-optimal algorithm. The greedy algorithm, formally studied by [20] in this context, is a natural contender. It is almost linear time, and easy to implement. However, [20] showed that such greedy algorithms can be really inaccurate. Moreover, even the greedy algorithm requires linear space and given the size of our data-sets, that can also become a constraint.
However, if our input instances admit a small sized covering then we show that the greedy algorithm run on a large enough random sample of data, i.e., Algorithm 1 for a large enough choice of $m$, performs reasonably well. Our main contribution here is Theorem 5.16. Theorem 5.16 states that the greedy algorithm on a small fraction random sample of an $\Theta(n)$ point data-set can be used to approximate the $1$-Wasserstein distance with a poly-logarithmic factor approximation.
So, the question arises: What precisely does a small covering assumption above mean and what for kind of natural problems does Algorithm 1 scales efficiently without deterioration in the approximation guarantee?
The metric entropy of a point set (see Definition 5.10) is the minimum number of balls of a given radius required to cover the point set. So, in high dimensions, data-sets with low metric entropy can be characterized from their values on (relatively) small balls spread through space. This is indeed the case with DEL data-sets like MNK1-MNK2 in our drug discovery application. The combinatorial synthesis process utilized in DELs often results in local chemical similarity among compounds that share common building blocks. Since similar molecules likely have the same binding behavior, synthesized molecules form a small ball around a parent molecule in the molecule fingerprint space. Therefore, molecule binding vs non-binding data-sets likely have low metric entropy. It turns out that for training data-sets with low metric entropy the greedy algorithm of [20] performs provably well (cf. Theorem 5.13).
The proof of Theorem 5.16 relies on Theorems 5.13 and 5.15.
Theorem 5.13 show that greedy minimum weight matching on bipartite graphs for vertex sets with low metric entropy has a much better (poly-logarithmic) approximation guarantee in our case, as opposed to the polynomial approximation guarantee from [20]. This argument, especially the connection between covering and matching in Lemma 5.12 may be of independent interest.
Theorem 5.15 essentially shows that random samples on data-sets with bounded metric entropy preserves $1$-Wasserstein distances. This allows us to work with small samples of large data-sets. For the proof, we need large deviation bounds for the $1$-Wasserstein distance between the theoretical distribution and its empirical distribution. Such results have been explored previously with tight Sanov’s theorem type bounds in low dimensional spaces (see for example [4]), but in high dimensions, we need the assumption of low metric entropy for the same results to go through (see chapter 6 in [9]). Coincidentally, that is precisely our assumption in Theorem 5.13!
5 Theorems and Proofs
5.1 Bounding $1$-Wasserstein distance suffices
In this section, in Theorem 5.4, we show that the Wasserstein distance between two measures, corresponding in the sum of squares loss function, upper bounds the total variation distance between the invariant measures underlying the stochastic gradient descent (SGD) algorithms.
Assumption 5.1.
We assume that our input consists of $2n$ points among the vertices of the hypercube: $Q_{d(n)}:=\{0,1\}^{d(n)}$, where $d(n)$ is $\Theta(\log n)$.
Restricting the state-space to the hypercube is fairly standard in algorithms literature. For the drug discovery example, the state-space is just the binary molecule fingerprint vectors i.e., roughly $d(n)=2000$. Note that assuming training and test data-sets of equal size is more for clarity of presentation, as one can add dummy example points of 00 weight, if needed.
Let $X\times Y$ denote the usual space of labeled examples i.e, in our case $X\subseteq\{0,1\}^{d(n)}$ is the set of feature values and $Y:=\{0,1\}$ is the set of labels. Our object of interest in this section is a neural network with smooth bounded activation functions. Let $y=f(w,x)$ denote the abstraction of our neural network, where $w$ denotes the real valued vector of weights. For a depth $p$ neural-net with polynomial activation functions of degree $q$, $f(w,x)$ is a polynomial in $x$ with degree at most $pq$.
Let $\ell(\cdot)$ denote the loss function, which we will assume to be the sum of square loss, for the sake of concreteness. The ideas easily extend to any low degree loss function. The training loss can be written as:
$\ell_{w}(\mathbb{P}_{\mathcal{S}}):=\mathbb{E}_{(x,y)\sim\mathbb{P}_{\mathcal{S}}}[(y-f(w,x))^{2}].$ (1)
Recall that (see for example [6]), a stochastic gradient descent algorithm with loss function $\ell$ can be abstracted as the Itó diffusion in the limit of small step size:
$\mathrm{d}w_{\mathcal{S}}(t)=\nabla_{w}\ell_{w}(\mathbb{P}_{\mathcal{S}})\mathrm{dt}+\sigma_{\mathcal{S}}\mathrm{dB(t)},$ (2)
where $\nabla_{w}$ denotes gradient with respect to $w$, $B(t)$ denotes Standard Brownian Motion in $|w|$-dimensions and the matrix $\sigma_{\mathcal{S}}$ depends upon the variance of the loss function for the mini-batch, mini-batch size and the learning rate.
Assumption 5.2.
We assume that the diffusion corresponds to an uniformly elliptic generator, since that ensures the existence of a unique limiting (invariant) measure [2]. Furthermore, we assume $\sigma_{\mathcal{S}}$ is isotropic i.e, it’s a scalar multiple of the identity $\sigma\cdot\mathrm{Id}$ and that $\sigma_{\mathcal{S}}=\sigma_{\mathcal{T}}$, in Theorem 5.4.
We relax the isotropy assumption somewhat in a corollary (see supplement). Finally, for our situation of interest, i.e., $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ large, we make a covariate shift type assumption.
Assumption 5.3.
We assume that $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})=\Omega(1)$. Furthermore, $f$ and $y$ are bounded, say $y,f\in[0,1]$ and
${\scriptstyle\left|\mathbb{E}_{y\sim\mathbb{P}_{\mathcal{S}}(\cdot|x)}[(y-f(w,x))^{2}]-\mathbb{E}_{y\sim\mathbb{P}_{\mathcal{T}}(\cdot|x)}[(y-f(w,x))^{2}]\right|=O(1).}$ (3)
Essentially, it says the data-sets have similar average loss in the same neighborhood for a given set of weights.
Theorem 5.4.
Suppose we train two neural networks, under the assumptions 5.2 and 5.3 above, on different input distributions, $\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}$, using the stochastic gradient descent (SGD) algorithm. Then, the total variation distance between their invariant measures can be bounded by $O(W_{1}(\mathbb{P}_{\mathcal{T}},\mathbb{P}_{\mathcal{S}}))$, in the limit as SGD step-size goes to 00.
Proof deferred to supplement.
Remark 5.5.
For dimension $d(n)$ large, the Levy-Prokhorov distance $L(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ between two distributions can be $\omega(1)$ times the Wasserstein distance $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$, so a Levy-Prokhorov metric based algorithm and guarantee can be weaker than the above.
5.2 The Greedy Algorithm
5.2.1 Reduction to bipartite matching
So far, we have established that $1$-Wasserstein metric is a sufficient topology to work with. This leads to the problem of computing the $1$-Wasserstein distance on two large datasets. That problem is equivalent to the minimum weight bipartite matching problem. In particular, we have the following lemma from [23].
Theorem 5.6.
[23] Given an instance of the optimal transport problem with supply and demands on two sets of points $(R,B)$, i.e, equivalently the $1$-Wasserstein distance computation problem in our case; we can construct an instance of the minimum weight bipartite matching problem such that solving the latter up to an approximation factor $\alpha$ will solve the former up to the same approximation factor $\alpha$.
The Algorithm ScaledGreedyWeight (below) carries out the reduction in Theorem 5.6 and calls GreedyMatch which matches two multisets embedded in a metric space using greedy algorithm on the edge lengths.
Notation: Scaling a discrete probability distribution $\mathbb{P}$ up by an integer factor of $C$ leads to a numerical rounding error of $\frac{1}{C\min\{\mathbb{P}\}}$, where ${}\min\{\mathbb{P}\}$ denotes the minimum positive value of density $\mathbb{P}$. Assume that we pick a large enough constant $C$ below, so that we can ignore the rounding error for the purposes of Theorem 5.13.
Algorithm 2 ScaledGreedyReweight (scale distributions and call bipartite matching)
1:Input: Two probability distributions $\mathbb{P}_{B},\mathbb{P}_{R}$ supported on $B,R\subset Q_{d}$, and a tilt factor $\alpha\in(0,1)$.
2:Output: Probability distribution $\mathbb{P}_{B}^{\prime}$ supported on $B$. $\mathbb{P}_{B}^{\prime}$ is close to $\alpha\mathbb{P}_{R}+(1-\alpha)\mathbb{P}_{B}$ in $W_{1}$, under assumptions of Theorem 5.13.
3:for $r\in R$ do
4:     $\mathrm{Supply}(r)\leftarrow C\cdot\alpha\mathbb{P}_{R}(r)$
5:for $b\in B$ do
6:     $\mathrm{Demand(b)}\leftarrow C-C\cdot(1-\alpha)\mathbb{P}_{B}(r)$
7:     if $\mathrm{Demand}(b)<0$ then
8:         $\mathrm{Demand(b)}\leftarrow 0$
9:Create multi-set $B^{\prime},R^{\prime}$ with multiplicities of each element being equal to their Demand and Supply respectively.
10:Use $\mathrm{GreedyMatch}(R^{\prime},B^{\prime})$ to compute the met (matched) demands, i.e., the extent to which the demands of $B$ that are actually fulfilled by $R$.
11:Normalize the weights of met demands to obtain a probability distribution $\mathbb{P}_{B}^{\prime}$ supported on $B$.
12:return $\mathbb{P}_{B}^{\prime}$.
5.2.2 Greedy algorithm and metric entropy
Recall that the data-points are set in the $d$-dimensional hypercube $Q_{d}$ with $\ell_{1}$ metric, where $d=\log^{O(1)}n$. The minimum weight bipartite matching problem is known to be harder than its non bipartite version. For example, the greedy algorithm is known to have a lower bound of $\Theta(n^{\log_{2}3/2})$ [20] for the bipartite version with $n$ data-set $\mathcal{T}$ and $n$ data-set $\mathcal{S}$ vertices. As an aside, a variant of the greedy algorithm (the hyper-greedy algorithm) provides a $\log n$ approximation in the non-bipartite case i.e, when any vertex can be matched to any other vertex. For the bipartite case, we obtain better approximation guarantees via the greedy algorithm, assuming small metric entropy of the input point sets.
Definition 5.7.
Given a perfect matching $M$ over a subset of vertices $C$ in a graph $G$, an alternating cycle $\gamma$ is a cycle in $G$ such that each alternate edge in the cycle belongs to $M$. Note that any such $M$ corresponds to a set of vertex disjoint alternating cycles.
In particular, Reingold and Tarjan [20] essentially show the following theorem.
Theorem 5.8.
[20] Given a set of $n$ data-set $\mathcal{T}$ and data-set $n$ $\mathcal{S}$ points in a metric space, the greedy algorithm returns a matching with weight that is within a factor of $\gamma^{\log_{2}\frac{3}{2}}$ of the minimum weight matching, where $\gamma$ is the length of the longest alternating cycle in the set (which can be $\Theta(n)$ for $Q_{\log n}$).
In order to improve upon their guarantee, we will exploit the following assumption for our input instance.
Assumption 5.9.
We assume that all input $\mathcal{T}$ and $\mathcal{S}$ points can be covered by $\eta$ balls of radius $\zeta$ lying within $Q_{d}$. We call such an input instance $(\eta,\zeta)$-bounded. The parameters $\eta$ and $\zeta$ will determine the approximation guarantee of our algorithm.
Definition 5.10.
Given a metric space, say $(Q,d)$ and $E\subset Q$,, the metric entropy $N^{\mathrm{ent}}_{r}(E)$ is the largest number of points $\{x_{1},\dots,x_{n}\}$ one can find in $E$ that are $r$-separated, i.e., $d(x_{i},x_{j})\geq r$ for all $i\neq j$.
Definition 5.11.
Given a metric space, say $(Q,d)$ and $E\subset Q$, the (external) covering number $N^{\mathrm{cov}}_{r}(E)$ is the fewest number of points $\{x_{1},\dots,x_{n}\in Q\}$ such that the $d$-balls $\{B(x_{1},r),\dots,B(x_{n},r)\}$ cover $E$.
Lemma 5.12 (Structural Lemma).
For an alternating cycle $\gamma$ induced by the greedy matching, if the weight of edges in the alternating cycle coming from the greedy matching is at least $\alpha$ times the weight of edges in the alternating cycle coming from the minimum weight matching then the metric entropy of $\gamma$ is large i.e, more precisely,
$\left(N^{\mathrm{ent}}_{\alpha/2}(\gamma)\cdot\frac{2d-\alpha}{\alpha}\right)^{\log_{2}3/2}\geq\alpha.$ (4)
Proof deferred to supplement. Note that an approximation factor of $d$ is trivial on $Q_{d}$ or on any set with $d_{\min}=1$ and $d_{\max}=d(n)$. The following corollary shows that the above indeed helps to improve upon the trivial bound for appropriately bounded instances.
Corollary 5.12.1.
Lemma 5.12 implies that the greedy algorithm achieves an approximation factor of $o(d^{3/4})$ on a $(d^{3/4},d^{3/4})$-bounded instance.
Proof.
We know that $N^{\mathrm{cov}}_{r}(E)\geq N^{\mathrm{ent}}_{r}(E)$ (see for example [9]). Therefore, Lemma 5.12 implies
$\alpha\leq\left(N^{\mathrm{cov}}_{\alpha/2}(\gamma)\cdot\frac{2d-\alpha}{\alpha}\right)^{\log_{2}3/2}.$ (5)
For $\alpha=d^{3/4}$, the right side of Equation 5 is $d^{log_{2}{3/2}}$, while the left side is $d^{3/4}$. Since $\log_{2}3/2<3/4$ we have a contradiction. Therefore, $\alpha=o(d^{3/4}).$
Of course, as the metric entropy decreases, the approximation factor improves, see for example the theorem below.
Theorem 5.13.
For $\eta=O(d^{\frac{1}{\xi\log_{2}3/2}})$, ($\xi>1$), Lemma 5.12 implies that the greedy algorithm achieves an approximation factor of $\max\{2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{\xi(1+\log_{2}(3/2))}}\right)\}$ on a $(\eta,\zeta)$-bounded minimum weight matching instance.
Together with Theorem 5.6, Theorem 5.13 implies that the greedy algorithm obtains the approximation factor on a $(\eta,\zeta)$-bounded Wasserstein distance computation instance. Proof deferred to supplement.
5.2.3 Small random samples suffice
In this subsection, we show that if the metric entropy is small, and so is the spread (see Definition 5.14) of the underlying distribution, then the empirical distribution of a much (polynomially) smaller sample is close to the actual distribution, in the 1-Wasserstein metric, with high probability.
Definition 5.14.
Let $\mu$ be the uniform distribution supported on a subset of vertices $Q$ of $Q_{d(n)}$. The spread of $\mu$, $S(\mu)$, is defined as:
[…]
where $d(\cdot,\cdot)$ denotes the $\ell_{1}$ distance on $Q_{d(n)}$.
Note that the spread is positive and greater than $1$, for any distribution defined on the hypercube, since the minimum value of $d(\cdot,\cdot)$ is $1$. In general, $S(\mu)$ can be a function of $d(n)$.
Theorem 5.15.
For a $(\eta,\zeta)$ coverable point-set, with $m=\alpha(n)\left(\eta 2^{\zeta}\right)$ and $\alpha(n)\in(0,1)$, the 1-Wasserstein distance between the empirical distribution and the true distribution of data-sets with bounded metric entropy obeys the following Sanov type concentration bound:
${\scriptstyle\exists\ \alpha(n)\to 0,\ \lim_{n\to\infty}\frac{1}{\eta 2^{\zeta}}\ln\mathbb{P}\left(W_{1}(\hat{\mu}_{m},\mu)\geq\log\log n+o(1)S(\mu)\right)\leq-\Omega(1).}$ (7)
The proof is in the supplement. It closely follows the covering based proof of multidimensional Cramer’s theorem in its metric entropy version (exercise 6.2.19 in [9]) with two main differences: (1) we need the topology induced by the Wasserstein metric instead of the Levy metric, and (2) our space has large dimension, i.e., say $\log n$, which depends upon $n$ [^7] This is one reason why we can’t directly apply the result from [4]. The constants in the exponential in their theorems will depend on $n$, and it’s not immediately clear whether the dependence will lead to a non-trivial result.. The second point requires us to be more careful with the covering argument, and so we only prove a relatively weaker result, with the help of the transportation inequality from [5].
5.2.4 Efficiency guarantee: Greedy with random sampling
Theorems 5.155.13, and 5.6 imply the following efficiency guarantee about the greedy minimum weight bipartite matching algorithm (GreedyMatch) on a random sample, and therefore Algorithm 1 as well.
Theorem 5.16.
Suppose we are given two data-sets with $\mathcal{S}$ and $\mathcal{T}$ that are weighted according to distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$. If,
  1. 1.
    $\mathcal{S}\cup\mathcal{T}$ admits a small covering: an $(\eta,\zeta)$ covering with $\eta,\zeta=O(\log^{c}n)$ and $\eta=O(\log^{c}n)$ for some constant $c\leq\frac{1}{1+\log_{2}(3/2)}$; and
  2. 2.
    $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ are sufficiently far apart: $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})\geq\log\log n+o(1)\max\{S(\mathbb{P}_{\mathcal{S}}),S(\mathbb{P}_{\mathcal{T}})\}$
then the greedy algorithm achieves an approximation ratio of $\max\left(2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{1+\log_{2}(3/2)}}\right)\right)$ with probability $1-o(1)$, when computed on a small random sample of $r(n)$ fraction of data-points and $r(n)\to 0$.
More succinctly, Theorem 5.16 states that the greedy algorithm on a small random sample can be used to approximate $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ on our data-sets $\mathcal{S}$ and $\mathcal{T}$, as long as the data-sets admit a small size covering using balls of small radius, and the two training weight distributions $\mathbb{P}_{\mathcal{S}}$ and $\mathbb{P}_{\mathcal{T}}$ are sufficiently different, which is the interesting case.
6 Acknowledgements
The author is grateful to Wen Torng, JW Feng, Jin Xu and Partick Riley for their help and advise with this paper.